# Fixing the Double Penalty in Data-Driven Weather Forecasting Through a Modified Spherical Harmonic Loss Function

### v1

@@ 1 Introduction
Beginning in 2023, the release of data-driven atmospheric forecasting models powered by deep neural network architectures began a revolution in medium-range weather forecasting, with some commenters (Bauer, 2024) anticipating that data-driven forecasting will soon supplant traditional numerical weather prediction (NWP) systems in all operational contexts. GraphCast (Lam et al., 2023), FourCastNet (Kurth et al., 2023), and Pangu-Weather (Bi et al., 2023) demonstrated forecast skill superior to that of the high-resolution forecast system (IFS) of the European Centre for Medium Range Weather Forecasts (ECMWF) at lead times (forecast lengths) up to 10 days. Since the publication of these models, the field has been joined by many others, including the Artificial Intelligence Forecasting System (AIFS) developed by ECMWF itself (Lang et al., 2024a).
From the standpoint of machine learning, atmospheric forecasting is a large-scale generative problem comparable to predicting the next frame of a video. As a typical example, the version of the GraphCast model deployed experimentally by the National Oceanic and Atmospheric Administration (NOAA) (NOAA, 2024) predicts the 6-hour forecast for six atmospheric variables at each of 13 vertical levels plus five surface variables, on a ¼° latitude/longitude grid, for about 86 million output degrees of freedom in aggregate. GraphCast takes two time-levels as input, so the input for this model has about 170 million degrees of freedom.
These first-generation data-driven weather models generally act as deterministic forecast systems, where each unique initial condition is mapped to a single forecast and verified against a “ground truth” from a data analysis system. The ERA5 atmospheric reanalysis (Hersbach et al., 2020) of ECWMF is most often used as the source of initial and verifying data for these forecast systems owing to its high quality and consistent behaviour from 1979 to present.
@@ 1.1 The problem of forecast smoothing
Unfortunately, despite their forecast skill deterministic data-driven forecast systems are universally understood to produce overly-smooth forecasts. A typical example of this behaviour is shown in figure 1 where a 3.5-day prediction of winter storm Eunice by the 13-level, ¼° GraphCast model is too weak and overly smooth. This smoothing results in an under-prediction of localized extreme events, and it makes the model less suitable for downstream tasks such as spectral nudging (Husain et al., 2024) and data assimilation (Slivinski et al., 2024).
The conventional wisdom is that this smoothing is something that can be fixed in the context of an ensemble forecasting system, which produces realizations from the space of potential future forecasts. GenCast (Price et al., 2025) and AIFS-CRPS (Lang et al., 2024b) directly produce a stochastic ¼° forecast given initial values and a source of random noise. SEEDS (Li et al., 2024) and ArchesWeatherGen (Couairon et al., 2024) are examples of models that predict variations around an ensemble mean, using the generative step to “fill in the blanks” around a smooth baseline. Of these examples, all but AIFS-CRPS use a diffusion technique with mean squared error (MSE) used as the de-noising loss function, while AIFS-CRPS instead uses the continuous ranked probability score (CRPS, Gneiting and Raftery (2007)) as its loss function to directly optimize the spread/error relationship of its produced ensemble.
@@
Lagerquist and Ebert-Uphoff (2022) also develops a variety of loss functions based on the same spatial methods (such as filtering and max-pooling) to verify forecasts of convective events like thunderstorms in evaluation of high-resolution, limited-area models.
@@ 2.1 Smoothing is optimal under mean squared error
When MSE is used as the loss function when training a data-driven model, the double penalty problem is more than annoyance: it encourages the model to generate unrealistically smooth predictions by reducing the amplitude of unpredictable scales. To show this quantitatively, consider the case of predicting a single variable. Let $Y=\operatorname{\mathcal{N}}(0,1)$ be the target, and let $X$ be the imperfect prediction of that target, modelled as a normal random variable with a standard deviation of $\sigma_{X}=\sqrt{\operatorname{\mathbb{E}}(X^{2})}$ and correlation coefficient of $\rho=\operatorname{\mathbb{E}}(XY)/\sigma_{X}$, where $\operatorname{\mathbb{E}}(\cdot)$ is the expectation operator. Writing $X$ in terms of a correlated and an uncorrelated component gives:
@@ 2.2 Spectral separation of the mean squared error
Figure 2: Amplitude ratio (solid) and coherence (dashed) for the spherical harmonic modes of temperature at 850hPa for 1° GraphCast during the training process. At top, values for 6h lead time during the single-step pre-training phase and at bottom, values for 6h–72h during the forecast rollout (batches 300,000–311,000, incrementing one step every 1,000 batches).
@@ 2.3 Spectrally adjusted mean squared error
Table 1: Fine-tuning curriculum for the ¼°/13-level version of GraphCast trained for this study, including the peak and terminal learning rates (LR) of the cosine annealing schedule used at each stage. The batch size was 8 throughout, and each stage had a warm-up period of 128 batches.
@@ 2.4 Fine-tuning methodology
The other training hyperparameters, including AdamW (Loshchilov and Hutter, 2019) hyperparameter settings and per-variable, per-level loss weightings were identical to those described by Lam et al. (2023).
@@ 3.1 Effective resolution
Conventional NWP models based on physics modelling are widely understood to have an effective resolution that is coarser than the model’s native grid resolution. Limits to effective resolution come from the limited fidelity of spatial or temporal discretization, from artificial diffusion or damping used to stabilize a model, and from sub-grid processes (such as turbulence) that must be imperfectly estimated rather than directly modelled. A model behaves unrealistically at scales finer than its effective resolution, typically providing insufficient variability and too-smooth solutions.
@@
The control model significantly smooths fine scales even after a single 6-hour forecast step, and that smoothing increases with the forecast lead time. If we somewhat arbitrarily draw the line of effective resolution at the point where the model has lost 25% of the per-wavenumber energy (corresponding to a ratio of power spectral densities of $0.75$ or an amplitude ratio of $\sqrt{0.75}$), the 5-day predictions of the control model reach that cutoff at wavenumber 32, corresponding to oscillations with a wavelength of about 1250km. The exact value of this effective resolution is not very sensitive to the cutoff ratio.
@@
For illustration, appendix B.1 shows amplitude spectra for select variables at various lead times, without normalizing by the spectral magnitude of the ground truth.
@@ 3.2 Lagged ensemble verification
Figure 5 shows the evolution of these statistics versus lead time for a selection of variables and levels, and more detailed evaluation of CRPS and eRMSE are shown in figures 10 and 11. AMSE-based fine tuning results in consistent improvements to the CRPS while the eRMSE sees little change, indicating that the fine tuning process produces a better-calibrated (more dispersive) ensemble without degrading overall predictive performance.
@@ 3.3 Hurricane prediction and extreme weather
Data-driven NWP models like GraphCast improve predictions of hurricane tracks relative to conventional NWP models. Since storms are guided by large-scale “steering flows” that have natural scales of thousands of kilometers, these predictions of storm position are relatively unaffected by the models’ limited effective resolutions but benefit from improvements in large-scale forecast skill. However, cyclones themselves are comparatively small, and predictions of the storm intensity are significantly affected by MSE-induced smoothing.
@@ 4 Discussion
This is not a property inherent to data-driven NWP. The alternate loss function based on (6) uses a spectral transform to separate the loss attributable to amplitude error from that attributable to decorrelation, encouraging the model to reproduce a realistic spectrum even if it can’t make an accurate prediction. When applied to the ¼°, 13-level version of GraphCast as an abbreviated fine-tuning step, we recover a model that has a much finer effective resolution, has improved CRPS-based verification in a lagged ensemble setting, and fixes the weak intensity bias in the prediction of tropical cyclones.
When fine-tuned autoregressively over multiple forecast steps, the model suffers from a small amount of smoothing at moderate scales. We speculate that this is because such autoregressive training has two objectives: forecasts are asked both to be accurate (and thus sharp, per (6)) and to be good initial conditions for the next forecast step. This latter goal is implicit, and it is not directly affected by the loss function used in training. Future work will consider the use of a replay buffer in training (like that of Chen et al. (2023)) to see if long-range forecast skill might be retained with even better prediction of amplitudes.
@@ 4.1 Alternative grids
In these cases, we speculate that much of the same benefit could be attained through multi-resolution decompositions such as wavelet lifting (Sweldens and Schröder, 2000). If suitable equivalents to scale-dependent variance and correlation could be defined, something like (6) could “morally” work even though it would lose an exact correspondence to an unmodified mean square error.
4.2 Applications to other fields
@@
The method can be mechanically applied whenever a spectral decomposition is possible, but additional value is only likely when a sub-aggregation of that spectrum is meaningful. This is most obviously possible in other areas of fluid dynamics, particularly the modelling of turbulent flows. Turbulent eddies are intermittent and their locations unpredictable, but they’re still present at predictable ratios.
@@ 4.3 Applications to ensemble modelling
Code availability
The GraphCast training code used in this work will be made available upon full publication of this paper. In the meantime, the corresponding author would welcome any inquiries.
Acknowledgements
The authors are grateful to Charlie Hébert-Pinard, Vikram Khade, and Hugo Vandenbroucke-Menu of the Implementation & Operational Services and Scientific Application Development sections of the Canadian Centre for Meteorological and Environmental Prediction for the retrained 1° GraphCast model shown in figure 2. The insights from intermediate stages of training were of great help in developing this work.
@@ Appendix A Relationship to maximum likelihood estimation
If we elevate this property from a fortunate coincidence to a simplifying assumption, we can treat the set of modes corresponding to a particular total wavenumber as random variables and apply the machinery of ensemble verification to individual, deterministic forecasts. The goal of producing realistic forecasts despite limited predictability is conceptually similar to the goal of maximum-likelihood estimation, so we consider here the effect of Kullback-Leibler (KL) divergence minimization. In the meteorology literature, the KL divergence is named the continuous ignorance score [Tödter and Ahrens, 2012], and it is sometimes used for ensemble verification.
@@ Appendix B Supplemental verification
B.1 Spectra
@@
B.2 Quantile/quantile plots
@@
Quantile-quantile plots show a joint cumulative density function, and we use them here to evaluate the overall realism of the forecasts produced by the control and AMSE AR12 models independently of the forecast skill. In figures 8 and 9, the x-location of each point is the labelled percentile of North American weather station observations for Northern Hemisphere winter and summer periods. The y-location of each point is the corresponding percentile for the HRES analysis or the 5-day forecasts produced by the control and AMSE AR12 models, interpolated to the station locations. For example, in the left panel of figure 8, the 98th percentile corresponds to an observed wind speed of about 11.5m/s, but the 98th percentile of the HRES analysis was about 10m/s.
@@
B.3 Details of the lagged ensemble verification
Brenowitz et al. [2024] uses several metrics to evaluate the quality of the lagged ensembles. In this work, we use the fair CRPS score, the ensemble root mean squared error (eRMSE), and the (naïve) spread-error ratio (SER). The eRMSE and SER statistics are each derived from their squared versions (ensemble mean squared error and variance/MSE ratio), which are evaluated pointwise and integrated over the grid. For an ensemble of $N_{e}$ members ($x_{1\ldots N_{e}}$) evaluated over $N_{date}$ forecasts with verifying analysis $y$, the corresponding formulas are:
$\operatorname{CRPS}(x,y)$ $=\frac{1}{N_{date}}\sum_{d=1}^{N_{date}}\sum_{i}\sum_{j}\mathrm{d}A(i,j)\left(\frac{1}{N_{e}}\sum_{k=1}^{N_{e}}|x_{k}(i,j)-y(i,j)|+\right.$

### v2

@@ Abstract
Machine Learning, ICML, Numerical Weather Prediction, Spherical Harmonics
@@ 1 Introduction
The models developed in Weyn et al. (2020) and Keisler (2022) suggested that deep neural networks might “solve” the problem of medium-range weather forecasting with data-driven machine learning models. In 2023, the release of GraphCast (Lam et al., 2023), FourCastNet (Kurth et al., 2023), and Pangu-Weather (Bi et al., 2023) demonstrated forecast skill that met or surpassed that of the high-resolution forecast system (IFS) of the European Centre for Medium Range Weather Forecasts (ECMWF) at lead times (forecast lengths) up to 10 days, and some commenters (Bauer, 2024) anticipated that data-driven forecasting would soon supplant traditional numerical weather prediction (NWP) in all operational contexts. Since the publication of these models, the field has been joined by many others, including the Artificial Intelligence Forecasting System (AIFS) developed by ECMWF itself (Lang et al., 2024a).
From the standpoint of machine learning, atmospheric forecasting is a large-scale generative problem comparable to predicting the next frame of a video. As a typical example, the version of the GraphCast model deployed experimentally by the National Oceanic and Atmospheric Administration (NOAA) (Sadeghi Tabas et al., 2025; NOAA, 2024) predicts the 6-hour forecast for six atmospheric variables at each of 13 vertical levels plus five surface variables, on a ¼° latitude/longitude grid, for about 86 million output degrees of freedom in aggregate. GraphCast takes two time-levels as input, so the input for this model has about 170 million degrees of freedom.
These first-generation data-driven weather models generally act as deterministic forecast systems, where each unique initial condition is mapped to a single forecast and verified against a “ground truth” from a data analysis system. The ERA5 atmospheric reanalysis (Hersbach et al., 2020) of ECMWF is most often used as the source of initial and verifying data for these forecast systems owing to its high quality and consistent behaviour from 1979 to present.
@@ 1.1 The Problem of Forecast Smoothing
Despite their overall forecast skill, deterministic data-driven forecast systems are universally understood to produce overly-smooth forecasts. A typical example of this behaviour is shown in figure 1 where a 3.5-day prediction of winter storm Eunice by the 13-level, ¼° GraphCast model is too weak and overly smooth. This smoothing results in an under-prediction of localized extreme events, and it makes the model less suitable for downstream tasks such as spectral nudging (Husain et al., 2024) and data assimilation (Slivinski et al., 2025).
This smoothing is most-discussed in relation to the prediction of gridded, global weather fields, but it is still present in models that have radically different architectures. Allen et al. (2025) develops a model that operates directly in observation space without an underlying grid that still produces smooth forecasts of the future, and Han et al. (2024) shows diminished forecast activity (a bulk measure related to blurring) at longer lead times for a local-area model despite a nominal kilometer-scale resolution.
The conventional wisdom is that this smoothing is something that can be fixed in the context of an ensemble forecasting system, which produces realizations from the space of potential future forecasts. GenCast (Price et al., 2025) and AIFS-CRPS (Lang et al., 2024b) directly produce a stochastic ¼° forecast given initial values and a source of random noise. SEEDS (Li et al., 2024) and ArchesWeatherGen (Couairon et al., 2024) are examples of models that predict variations around an ensemble mean, using the generative step to “fill in the blanks” around a smooth baseline. Lippe et al. (2023) approaches this problem from a more general partial differential equation framework, and it develops a diffusion method that iteratively refines finer scales.
Of these examples, all but AIFS-CRPS use a diffusion technique with mean squared error (MSE) used as the de-noising loss function, while AIFS-CRPS instead uses the continuous ranked probability score (CRPS, Gneiting & Raftery (2007)) as its loss function to directly optimize the spread/error relationship of its produced ensemble.
@@
Lagerquist & Ebert-Uphoff (2022) also develops a variety of loss functions based on the same spatial methods (such as filtering and max-pooling) to verify forecasts of convective events like thunderstorms in evaluation of high-resolution, limited-area models.
@@ 2.1 Smoothing Is Optimal Under Mean Squared Error
When MSE is used as the loss function to train a data-driven model, the double penalty problem is more than annoyance: it encourages the model to generate unrealistically smooth predictions by reducing the amplitude of unpredictable scales. To show this quantitatively, consider the case of predicting a single variable. Let $Y=\operatorname{\mathcal{N}}(0,1)$ be the target, and let $X$ be the imperfect prediction of that target, modelled as a normal random variable with a standard deviation of $\sigma_{X}=\sqrt{\operatorname{\mathbb{E}}(X^{2})}$ and correlation coefficient of $\rho=\operatorname{\mathbb{E}}(XY)/\sigma_{X}$, where $\operatorname{\mathbb{E}}(\cdot)$ is the expectation operator. Writing $X$ in terms of a correlated and an uncorrelated component gives:
@@ 2.2 Spectral Separation of the Mean Squared Error
Figure 2: Amplitude ratio (solid) and coherence (dashed) for the spherical harmonic mode with total wavenumber 100 for temperature at 850hPa during the training of a 1° version of the GraphCast model with an MSE loss function. At top, values for 6h lead time during the single-step pre-training phase and at bottom, values for 6h–72h during the forecast rollout (batches 300,000–311,000, incrementing one step every 1,000 batches).
@@ 2.3 Spectrally Adjusted Mean Squared Error
Table 1: Fine-tuning curriculum for the ¼°/13-level version of GraphCast trained for this study, including the peak and terminal learning rates (LR) of the cosine annealing schedule used at each stage. The batch size was 8 throughout, and each stage had a warm-up period of 64 batches.
@@ 2.4 Fine-Tuning Methodology
The other training hyperparameters, including AdamW (Loshchilov & Hutter, 2019) hyperparameter settings and per-variable, per-level loss weightings were identical to those described by Lam et al. (2023).
@@ 3.1 Effective Resolution
Conventional, physics-based NWP models are widely understood to have an effective resolution that is coarser than the model’s native grid resolution. Limits to effective resolution come from the limited fidelity of spatial or temporal discretization, from artificial diffusion or damping used to stabilize a model, and from sub-grid processes (such as turbulence) that must be imperfectly estimated rather than directly modelled. A model behaves unrealistically at scales finer than its effective resolution, typically providing insufficient variability and too-smooth solutions.
@@
The control model significantly smooths fine scales even after a single 6-hour forecast step, and that smoothing increases with the forecast lead time. If we somewhat arbitrarily draw the line of effective resolution at the point where the model has lost 25% of the per-wavenumber energy (corresponding to a ratio of power spectral densities of $0.75$ or an amplitude ratio of $\sqrt{0.75}$), the 5-day predictions of the control model reach that cutoff at wavenumber 32, corresponding to oscillations with a wavelength of about 1250 km. Small changes in the target amplitude ratio will result in small changes to the derived effective resolution.
@@
For illustration, appendix B.2 shows amplitude spectra for select variables at various lead times, without normalizing by the spectral magnitude of the ground truth. Appendix B.5 discusses the effective resolution of the model when trained with either mean squared error or mean absolute (L1) error.
@@ 3.2 Lagged Ensemble Verification
Figure 4 shows the evolution of these statistics versus lead time for a selection of variables and levels, and more detailed evaluation of CRPS and eRMSE are shown in figures 11 and 12. The AMSE AR12 model shows consistent improvements to the CRPS while the eRMSE sees little change, indicating that the fine tuning process produces a better-calibrated (more dispersive) ensemble without degrading overall predictive performance.
@@ 3.3 Hurricane Prediction and Extreme Weather
Data-driven NWP models like GraphCast improve predictions of hurricane tracks relative to conventional NWP models (see for example figure 3A of Lam et al. (2023)). Since storms are guided by large-scale “steering flows” that have natural scales of thousands of kilometers, these predictions of storm position are relatively unaffected by the models’ limited effective resolutions but benefit from improvements in large-scale forecast skill. However, cyclones themselves are comparatively small, and predictions of the storm intensity are significantly affected by MSE-induced smoothing.
@@ 4 Discussion & Limitations
This is not a property inherent to data-driven NWP. The alternate loss function based on (6) uses a spectral transform to separate the loss attributable to amplitude error from that attributable to decorrelation, encouraging the model to reproduce a realistic spectrum even if it can’t make an accurate prediction. When applied to the ¼°, 13-level version of GraphCast with an abbreviated fine-tuning process, we recover a model that has a much finer effective resolution, has improved CRPS-based verification in a lagged ensemble setting, and fixes the weak intensity bias in the prediction of tropical cyclones.
When fine-tuned autoregressively over multiple forecast steps, the model suffers from a small amount of smoothing at mesoscales (intermediate scales). We speculate that this is because such autoregressive training has two objectives: forecasts are asked both to be accurate (and thus sharp, per (6)) and to be good initial conditions for the next forecast step. This latter goal is implicit, and it is not directly affected by the loss function used in training. Future work will consider the use of a replay buffer in training (like that of Chen et al. (2023)) to see if long-range forecast skill might be retained with even better prediction of amplitudes.
Since the AMSE loss function (6) is zero if and only if the predicted field matches the ground truth, it may be useful throughout model training rather than just during a fine-tuning pass. However, a thorough test of this proposition would require a considerable computational budget, so it is left for future work. Use of the AMSE loss function throughout the training process might improve the coherence of fine-scale prediction by allowing the model to spend more of its training time “seeing” these modes, but on the other hand the coherence-dependent smoothing encouraged by the MSE loss function (figure 2) might act as an implicit regularization that smooths the model’s gradients and speeds up training overall.
4.1 Effective Resolution
The ultimate conclusion of this work is that the AMSE-based error measure improves the effective resolution of NWP weather models, but the phrase “effective resolution” must always be accompanied by the question, “effective at what?”
We chose to define an effective resolution based on smoothing of fine scales, since a model that simply doesn’t represent a scale cannot effectively model it. However, other definitions exist in the literature, and users of these models should keep their ultimate goals in mind. For example, Kent et al. (2014) studies various discretization schemes for numerical partial differential equations under both diffusion (smoothing) and dispersion (wave propagation) definitions of effective resolution.
@@ 4.2 Alternative Grids
In these cases, we think that the basic intuition behind (6) might still apply through other multiscale decompositions such as wavelet lifting (Sweldens & Schröder, 2000), provided suitable equivalents to scale-dependent variance and correlation could be found. The multiscale decomposition is critical in some form, however, since the method takes advantage of the approximate independence of scale-separated modes. Without such a decomposition (e.g. applying the adjustment of (6) globally, without the harmonic transform), the model might be able to “cover up” a lack of fine-scale variability by over-emphasizing coarser scales.
4.3 Applications to Other Domains
@@
The method can be mechanically applied whenever a spectral decomposition is possible, but additional value is only likely when a sub-aggregation of that spectrum is meaningful. This is most obviously possible in other areas of fluid dynamics, particularly the modelling of turbulent flows. In that domain, Chakraborty et al. (2025) developed a binned spectral loss function (on a planar domain) that is reminiscent of the amplitude-only component of (6), but it discards the phase information. We are optimistic that integrating the spectral correlation along the lines of AMSE will make such models more robust.
@@ 4.4 Applications to Ensemble Modelling
Code and Data Availability
An implementation of the AMSE error function and the code used to train GraphCast for this work are available at https://github.com/csubich/graphcast/tree/amse under the Apache 2.0 license. The fine-tuned checkpoints produced for this study are available at https://huggingface.co/csubich/graphcast_amse under the CC-BY-ND-SA 4.0 license, as derivative works of the DeepMind “graphcast-operational” checkpoint.
Acknowledgements
The authors would like to thank Charlie Hébert-Pinard, Vikram Khade, and Hugo Vandenbroucke-Menu of the Canadian Centre for Meteorological and Environmental Prediction for access to the 1°-trained version of GraphCast used to produce the results of figure 2.
@@ Appendix A Relationship to Maximum Likelihood Estimation
If we elevate this property from a fortunate coincidence to a simplifying assumption, we can treat the set of modes corresponding to a particular total wavenumber as random variables and apply the machinery of ensemble verification to individual, deterministic forecasts. The goal of producing realistic forecasts despite limited predictability is conceptually similar to the goal of maximum-likelihood estimation, so we consider here the effect of Kullback-Leibler (KL) divergence minimization. In the meteorology literature, the KL divergence is named the continuous ignorance score (Tödter & Ahrens, 2012), and it is sometimes used for ensemble verification.
@@ Appendix B Supplemental Verification
Figure 7: Visualization of high-pass filtered forecast and analysis fields for the forecast shown in 1.
B.1 Visualization
Figure 7 visualizes the high-wavenumber components of a sample forecast matching the winter storm Eunice prediction shown in figure 1. The applied filter fourth-order in spherical harmonic space, with the functional form:
$\textrm{HPF}(k)=1-\frac{k_{0}^{4}}{k_{0}^{4}+k^{4}},$ (13)
where $k$ is the total wavenumber and $k_{0}=50$ is the cutoff number, chosen to emphasize modes with length scales of 800 km and shorter. Overall, the predictions of the control and AMSE-trained models show very similar structures, but training with (6) as the loss function enhances the high-mode variability of the forecasts.
B.2 Spectra
@@
B.3 Quantile/Quantile Plots
@@
Quantile-quantile plots show a joint cumulative density function, and we use them here to evaluate the overall realism of the forecasts produced by the control and AMSE AR12 models independently of the forecast skill. In figures 9 and 10, the x-location of each point is the labelled percentile of North American weather station observations for Northern Hemisphere winter and summer periods. The y-location of each point is the corresponding percentile for the HRES analysis or the 5-day forecasts produced by the control and AMSE AR12 models, interpolated to the station locations. For example, in the left panel of figure 9, the 98th percentile corresponds to an observed wind speed of about 11.5m/s, but the 98th percentile of the HRES analysis was about 10m/s.
@@
B.4 Details of the Lagged Ensemble Verification
Brenowitz et al. (2025) uses several metrics to evaluate the quality of the lagged ensembles. In this work, we use the fair CRPS score, the ensemble root mean squared error (eRMSE), and the spread-error ratio (SER). The eRMSE statistic is derived from its squared version (ensemble mean squared error), evaluated pointwise and integrated over the grid. The SER statistic is the simple ratio of the integrated MSE and ensemble spread (unbiased estimate of variance), noting for emphasis that the ratio is taken after the grid-averaging. For an ensemble of $N_{e}$ members ($x_{1\ldots N_{e}}$) evaluated over $N_{date}$ forecasts with verifying analysis $y$, the corresponding formulas are:
$\operatorname{CRPS}(x,y)$ $=\frac{1}{N_{date}}\sum_{d=1}^{N_{date}}\sum_{i,j}\mathrm{d}A(i,j)\left(\frac{1}{N_{e}}\sum_{k=1}^{N_{e}}|x_{k}(i,j)-y(i,j)|+\right.$